OpenStudy (anonymous):

EASY MEDAL :P Would it be B or C? Which expression is equivalent to (73)−2? 1 over 7 times 7 times 7 times 7 times 7 times 7 7 1 over 7 negative 1 over 7 times 7 times 7 times 7 times 7 times 7

OpenStudy (anonymous):

um lol your question is soooo hard to read!

OpenStudy (anonymous):

let me screenshot it :)

OpenStudy (anonymous):

\[\frac{ 1 }{ 7 \times 7 \times 7 \times 7 \times 7 \times 7}\]

OpenStudy (anonymous):

what do you think it is?

OpenStudy (anonymous):

the expression looks like\[(73)^{-2}=\frac{ 1 }{ 73^2 }\]

OpenStudy (anonymous):

1/117649

OpenStudy (anonymous):

c